# Graphical lasso for extremes

### v2

@@ Appendix A Proof of propositions in Sections 2 and 3
From $\Sigma\cdot\mathbf{1}=\mathbf{0}$, we have that $\frac{1}{\sqrt{d}}\cdot\mathbf{1}$ is an eigenvector of $\Sigma$ with eigenvalue 00. Since $\Sigma$ is positive semidefinite with rank $d-1$. it follows that the rest of the eigenvalues $\lambda_{2}\leq\cdots\leq\lambda_{d}$ must be positive.
@@
$-\left(\hat{\Theta}^{*}\right)^{-1}+S^{*}+\gamma_{n}\hat{Z}=0,$ (B.1)
@@
$\hat{Z}_{ij}=\begin{cases}0&\text{ if }i=j,\\ \text{sign}(\hat{\Theta}^{*}_{ij}-c)&\text{ if }i\neq j\text{ and }\hat{\Theta}^{*}_{ij}\neq c,\\ \in[-1,1]&\text{ if }i\neq j\text{ and }\hat{\Theta}^{*}_{ij}=c.\end{cases}$
@@
$-\left(\tilde{\Theta}^{*}\right)^{-1}_{ij}+S^{*}_{ij}+\gamma_{n}\tilde{Z}_{ij}=0,\quad(i,j)\in E,$
@@
Then the pair $(\tilde{\Theta}^{*},\tilde{Z})$ satisfies the original KKT equation (B.1). What remains to be proved is that $\tilde{Z}$ also satisfies
@@
Now in the case where $|||\Sigma^{*}\Delta|||_{\infty}<1$, we can expand $J$ as
@@
as the approximation error. Here $R$ is defined regardless of whether $|||\Sigma^{*}\Delta|||_{\infty}<1$.
@@
  • $-\left(\Theta^{*}+\Delta\right)^{-1}_{E}+S^{*}_{E}+\gamma_{n}\tilde{Z}_{E}=0$;
@@
$\|F(\overline{\Delta}_{E})\|_{\infty}\leq|||(\Omega_{EE})^{-1}|||_{\infty}\left(\|R\|_{\infty}+\|R^{\prime}\|_{\infty}+\gamma_{n}\|\overline{\tilde{Z}}_{E}\|_{\infty}\right)\leq|||(\Omega_{EE})^{-1}|||_{\infty}\left(\|R\|_{\infty}+\|R^{\prime}\|_{\infty}+\gamma_{n}\right),$
@@
provided that $|||\Sigma^{*}\Delta|||_{\infty}<1$. To ensure this condition,
$|||\Sigma^{*}\Delta|||_{\infty}=|||\Sigma^{*}(\tilde{\Theta}-\Theta)|||_{\infty}\leq D|||\Sigma^{*}|||_{\infty}\cdot r_{n},$
where $D$ is the maximum degree in the graph. Therefore $|||\Sigma^{*}\Delta|||_{\infty}<1$ holds by requiring that
$D|||\Sigma^{*}|||_{\infty}\cdot r_{n}\leq C_{4}<1.$ (B.5)
The upper bound $C_{4}$ implies that $|||J^{T}|||_{\infty}\leq\frac{1}{1-C_{4}}.$
@@
$R_{ij}=\mathbf{e}_{i}^{T}(\Sigma^{*}\Delta)^{2}J\Sigma^{*}\mathbf{e}_{j}\leq\|\mathbf{e}_{i}^{T}(\Sigma^{*}\Delta)^{2}\|_{\infty}\|J\Sigma^{*}\mathbf{e}_{j}\|_{1}\leq\|(\Sigma^{*}\Delta)^{2}\|_{\infty}|||\Sigma^{*}J^{T}|||_{\infty}.$
@@
$\|R\|_{\infty}$$\leq$ $\|(\Sigma^{*}\Delta)^{2}\|_{\infty}|||\Sigma^{*}J^{T}|||_{\infty}$
$\leq$ $|||\Sigma^{*}\Delta|||_{\infty}\|\Sigma^{*}\Delta\|_{\infty}|||J^{T}|||_{\infty}|||\Sigma^{*}|||_{\infty}$
$\leq$ $D|||\Sigma^{*}|||_{\infty}\cdot r_{n}\cdot|||\Sigma^{*}|||_{\infty}r_{n}\cdot\frac{1}{1-C_{4}}\cdot|||\Sigma^{*}|||_{\infty}$
@@
$\|F(\Delta_{E})\|_{\infty}\leq|||(\Omega_{EE})^{-1}|||_{\infty}(C_{5}\cdot r^{2}_{n}+\delta_{n}+\gamma_{n})\leq r_{n},$
@@
$C_{5}\cdot r_{n}^{2}+\delta_{n}+\gamma_{n}\leq\frac{1}{|||(\Omega_{EE})^{-1}|||_{\infty}}\cdot r_{n}$ (B.6)
@@
$\|\overline{\tilde{Z}}_{E^{c}}\|_{\infty}\leq\frac{1}{\gamma_{n}}|||\Omega_{EE^{c}}(\Omega_{EE})^{-1}|||_{\infty}\left(\|R\|_{\infty}+\|R^{\prime}\|_{\infty}+\gamma_{n}\|\overline{\tilde{Z}}_{E}\|_{\infty}\right)+\frac{1}{\gamma_{n}}\|R\|_{\infty}+\frac{1}{\gamma_{n}}\|R^{\prime}\|_{\infty}.$
@@
$\|\overline{\tilde{Z}}_{E^{c}}\|_{\infty}$$\leq$ $\frac{1}{\gamma_{n}}|||\Omega_{EE^{c}}(\Omega_{EE})^{-1}|||_{\infty}\left(\delta_{n}+C_{5}\cdot r_{n}^{2}+\gamma_{n}\right)+\frac{1}{\gamma_{n}}(\delta_{n}+C_{5}\cdot r_{n}^{2})$
$=$ $|||\Omega_{EE^{c}}(\Omega_{EE})^{-1}|||_{\infty}+\frac{1}{\gamma_{n}}\left(|||\Omega_{EE^{c}}(\Omega_{EE})^{-1}|||_{\infty}+1\right)\left(\delta_{n}+C_{5}\cdot r_{n}^{2}\right).$
Since Condition (4.1) ensures that $|||\Omega_{EE^{c}}(\Omega_{EE})^{-1}|||_{\infty}<1-\alpha$, To satisfy $\|\overline{\tilde{Z}}_{E^{c}}\|_{\infty}\leq 1$, we only need to further require
@@
$r_{n}:=\frac{|||\left(\Omega_{EE}\right)^{-1}|||_{\infty}}{1-\alpha}\cdot\gamma_{n}.$
@@
$D|||\Sigma^{*}|||_{\infty}\frac{|||(\Omega_{EE})^{-1}|||_{\infty}}{1-\alpha}\cdot\gamma_{n}$$\leq$ $C_{4}$
$\delta_{n}+\frac{D}{1-C_{4}}|||\Sigma^{*}|||_{\infty}^{3}\cdot\left(\frac{|||(\Omega_{EE})^{-1}|||_{\infty}}{1-\alpha}\right)^{2}\cdot\gamma_{n}^{2}$$\leq$ $\frac{\alpha}{1-\alpha}\cdot\gamma_{n}.$
where $C_{5}$ is substituted by $\frac{D}{1-C_{4}}|||\Sigma^{*}|||_{\infty}^{3}$:
The lower bound for $\gamma_{n}$ in (4.2) ensures that $\delta_{n}\leq\epsilon\frac{\alpha}{1-\alpha}\gamma_{n}$ for some $10<\epsilon<$ 1, we thus need to require that
$\frac{D}{1-C_{4}}|||\Sigma^{*}|||_{\infty}^{3}\cdot\left(\frac{|||(\Omega_{EE})^{-1}|||_{\infty}}{1-\alpha}\right)^{2}\cdot\gamma_{n}\leq(1-\epsilon)\frac{\alpha}{1-\alpha},$
@@
$\gamma_{n}\leq\min\left\{\frac{C_{4}(1-\alpha)}{D|||\Sigma^{*}|||_{\infty}|||(\Omega_{EE})^{-1}|||_{\infty}},\frac{(1-C_{4})(1-\epsilon)\alpha(1-\alpha)}{D|||\Sigma^{*}|||_{\infty}^{3}|||(\Omega_{EE})^{-1}|||_{\infty}^{2}}\right\}.$
@@
$C_{4}=\frac{(1-\epsilon)\alpha}{(1-\epsilon)\alpha+|||\Sigma^{*}|||_{\infty}^{2}|||(\Omega_{EE})^{-1}|||_{\infty}}<1,$
@@
$\gamma_{n}\leq\frac{(1-\epsilon)\alpha(1-\alpha)}{D|||\Sigma^{*}|||_{\infty}|||(\Omega_{EE})^{-1}|||_{\infty}\left[(1-\epsilon)\alpha+|||\Sigma^{*}|||_{\infty}^{2}|||(\Omega_{EE})^{-1}|||_{\infty}\right]}.$
@@
$|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$ $=$ $\sup_{\mathbf{x}\in\mathbb{R}^{|E|}}\frac{\|\Omega_{E^{c}E}(\Omega_{EE})^{-1}\mathbf{x}\|_{\infty}}{\|\mathbf{x}\|_{\infty}}$
@@
$|||\Omega_{E^{c}E}(\Omega_{EE})^{-1}|||_{\infty}$$\leq$ $\left(|E|\cdot\left(\frac{\lambda_{\max}((\Omega^{2})_{EE})}{\lambda_{\min}((\Omega_{EE})^{2})}-1\right)\right)^{1/2}$
@@ E.1 Problem specification
$-\Theta^{-1}+S+\lambda\Gamma=\mathbf{0},$ (E.2)
@@ E.3 P-GLASSO algorithm
$-W+S+\lambda\Gamma=\mathbf{0}.$ (E.11)
@@
$-\mathbf{w}_{12}+\mathbf{s}_{12}+\lambda\bm{\gamma}_{12}=\mathbf{0}.$ (E.12)
@@ E.4 Pseudo code: P-GLASSO for extremes
$\mathbf{w}_{12}^{T}\bm{\theta}_{12}+w_{22}\cdot\theta_{22}$ $=$ $1$ (E.17)

### v3

@@ Appendix A Proof of propositions in Sections 2 and 3
From $\Sigma\cdot\mathbf{1}=\mathbf{0}$, we have that $\frac{1}{\sqrt{d}}\cdot\mathbf{1}$ is an eigenvector of $\Sigma$ with eigenvalue 0. Since $\Sigma$ is positive semidefinite with rank d1d-1. it follows that the rest of the eigenvalues $\lambda_{2}\leq\cdots\leq\lambda_{d}$ must be positive.
@@
The proof of Theorem 4.1 follows the same strategy as that of Theorem 1 of Ravikumar et al. (2011). The key difference arises from the shifted penalty target $c>0$ in the extreme graphical lasso, which affects the construction of the witness solution and requires a modified analysis of the continuity of the mapping $F$. We highlight these differences as they arise in the proof.
@@
$\left(\hat{\Theta}^{*}\right)^{-1}+S^{*}+\gamma_{n}\hat{Z}=0,$ (B.1)
@@
[…]
@@
$\left(\tilde{\Theta}^{*}\right)^{-1}_{ij}+S^{*}_{ij}+\gamma_{n}\tilde{Z}_{ij}=0,\quad(i,j)\in E,$
@@
[…]
@@
Now in the case where $|\Sigma^{*}\Delta|||_{\infty}<1$, we can expand $J$ as
@@
as the approximation error. Here $R$ is defined regardless of whether $|\Sigma^{*}\Delta|||_{\infty}<1$.
@@
  • $\left(\Theta^{*}+\Delta\right)^{-1}_{E}+S^{*}_{E}+\gamma_{n}\tilde{Z}_{E}=0$;
@@
[…]
@@
provided that $|\Sigma^{*}\Delta|||_{\infty}<1$. To ensure this condition,
[…]
where $D$ is the maximum degree in the graph. Therefore $|\Sigma^{*}\Delta|||_{\infty}<1$ holds by requiring that
[…]
The upper bound $C_{4}$ implies that $|J^{T}|||_{\infty}\leq\frac{1}{1-C_{4}}.$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
Since Condition (4.1) ensures that $|\Omega_{EE^{c}}(\Omega_{EE})^{-1}|||_{\infty}<1-\alpha$, To satisfy $\|\overline{\tilde{Z}}_{E^{c}}\|_{\infty}\leq 1$, we only need to further require
@@
[…]
@@
[…]
The lower bound for $\gamma_{n}$ in (4.2) ensures that $\delta_{n}\leq\epsilon\frac{\alpha}{1-\alpha}\gamma_{n}$ for some $0<\epsilon<1$, we thus need to require that
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ E.1 Problem specification
$\Theta^{-1}+S+\lambda\Gamma=\mathbf{0},$ (E.2)
@@ E.3 P-GLASSO algorithm
$W+S+\lambda\Gamma=\mathbf{0}.$ (E.11)
@@
$\mathbf{w}_{12}+\mathbf{s}_{12}+\lambda\bm{\gamma}_{12}=\mathbf{0}.$ (E.12)
@@ E.4 Pseudo code: P-GLASSO for extremes
$\mathbf{w}_{12}^{T}\bm{\theta}_{12}+w_{22}\cdot\theta_{22}$ $=$ $1.$ (E.17)